OpenStudy (anonymous):

if f(x)=2x+2 and g(x)=x-1 find f(g(4))

OpenStudy (anonymous):

g(4)=4-1=3 f(g(4)=f(3)=2*3+2=8

OpenStudy (anonymous):

the only asnwers available are f(g(4))=3, f(g(4))=9, f(g(4))=8, f(g(4))=10

OpenStudy (anonymous):

also how did you get it?

OpenStudy (bibby):

he just wrote it out. find g(4) plug that into f(x)

OpenStudy (anonymous):

ah okay thank you